A building in the shape of a cube has a width of 45ft a length of 45ft and a height of 45ft . An architect want to build a scale model of the building that is 30% smaller in volume . What is the height of the new building ? Round to the nearest hundredth
Answer by Theo(13342)

the easiest way to figure this out is as follows:

45^3 = 91125

91125 - .3 * 91125 = 91125 * (1 - .3) = 91125 * .7 = 63787.5

63787.5 ^ (1/3) = 39.95568008

that's the new measure for the length, width, and height.

your solution is that the height of the new building is 39.96 feet.